Shift Summary for owl shift, 01/04/2000


Subject: Shift Summary for owl shift, 01/04/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:37:32 CST


              Shift Summary for owl shift, 01/04/2000

On Shift: J. Shields, Bob T.

Hours of beam available at all intensities: 2.5
Hours of beam available at nominal intensity: 2.5
Hours of E799 data (special runs not included): 2.5
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:

Booster RF, MI power supply probs
========================================================================
Runs ended this shift:
Run 15388 -- BEAM_799
Run 15389 -- LASER_SCAN

Tapes brought to FCC this shift:
  (including tapes for online split)

========================================================================
Accesses during shift:

========================================================================
Overview of shift:

Tevatron down due to Booster RF, MI power supply probs, E799 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15388 00:00 we inherit control room just as previous shift
                        crew finishes converting the KAMI test
                        configuration back into good ole BEAM_799.

              ~01:15 Beam dies. Booster RF and MI power supply
                        probs. No estimate.

15389 03:08 beam is apparently nowhere in sight, so shift
                        crew seizes the initiative and begins a laser scan.

               03:30 estimate for beam is 3 hours. (sigh)

               03:53 the infamous Mr. Burns "Exxx-cell-ent!" voice
                       startles the shift crew out of their sleepy
                       stupor and prompts them to stop the laser scan.

15390 04:03 A new BEAM_799 run is initiated, so
                        that we're ready for beam when it returns.

               05:25 beam returns an hour ahead of the estimate!!

               06:40 scalar out of limits alarm, but they were
                        only off by a little.

               06:46 kumquat module 4X-1 is not reading out. Reloaded
                        kumquats, but-- alas-- to no avail!
                           Unfazed by their failure, the intrepid shift
                        crew smoothly shifts into plan B: Lucky Jason
                        Ladue is treated to an early morning wake up call.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:37:44 CST